3 is subtracted from the product of 4 and a number
Found 2 solutions by ikleyn, farohw:
Answer by ikleyn(52803)
Answer by farohw(175)


Let the product of 4 and a number be 4x then subtract 3 from it:

The equation is 4x - 3

If you have to solve it, first set the equation equal to zero:
4x - 3 = 0

4x = 3

4x/4 = 3/4

x = 3/4